### v1

@@ Abstract
We investigate the energy efficiency of a library designed for parallel computations with sparse matrices. The library leverages high-performance, energy-efficient Graphics Processing Unit (GPU) accelerators to enable large-scale scientific applications. Our primary development objective was to maximize parallel performance and scalability in solving sparse linear systems whose dimensions far exceed the memory capacity of a single node.
To this end, we devised methods that expose a high degree of parallelism while optimizing algorithmic implementations for efficient multi-GPU usage. Previous work has already demonstrated the library’s performance efficiency on large-scale systems comprising thousands of NVIDIA GPUs, achieving improvements over state-of-the-art solutions.
In this paper, we extend those results by providing energy profiles that address the growing sustainability requirements of modern HPC platforms. We present our methodology and tools for accurate runtime energy measurements of the library’s core components and discuss the findings. Our results confirm that optimizing GPU computations and minimizing data movement across memory and computing nodes reduces both time-to-solution and energy consumption. Moreover, we show that the library delivers substantial advantages over comparable software frameworks on standard benchmarks.
@@ 4.1 CPU and GPU Power
powerMonitor has been adapted from GPowerU to better meet our specific requirements, providing a simple, yet effective, solution for monitoring the power consumption of all GPUs on a compute node. Indeed, it can be used either as an external tool or integrated directly into the application code. In our experiments, we used it as an external tool. In this case, it must be launched before the application starts and stopped once execution is complete. The complete workflow for CPU and GPU power monitoring is shown in Figure 1. powerMonitor creates an output file for each device found on the node, containing a sequence of power samplings along with the corresponding timestamps. These samples can be used to reconstruct the power timeline of each device during the application’s execution. Figure 2 shows an example of a power–time curve.
@@ 5 Results and Comparisons
5.1 SpMV Operation
We begin by analyzing the performance behavior of the SpMV operation under both strong and weak scalability scenarios.For the single-GPU case, the problem size is set to $405^{3}$ DOFs for the 7-point stencil and $260^{3}$ DOFs for the 27-point stencil, corresponding to the largest dimensions that fully saturate the available GPU memory. All performance measurements reported in the figures represent averages over five independent runs, with each run consisting of 100 repetitions of the SpMV computation.
@@ 5.1 SpMV Operation
Finally, Figure 6 presents the dynamic energy consumption per DOFs. As anticipated from previous results, BootCMatchGX demonstrates significantly higher energy efficiency than Ginkgo—approximately twice as efficient. In both implementations, energy efficiency remains nearly constant with increasing numbers of GPUs under weak scalability, confirming the good scalability of both solutions. Under strong scalability, as expected, efficiency declines as the number of GPUs increases, due to the reduction of the computational workload per GPU.
@@
5.2 Un-preconditioned Conjugate Gradient solver
@@
For the single-GPU case, the problem size is set to $408^{3}$ DOFs with the 7-point stencil and $265^{3}$ DOFs with the 27-point stencil; AmgX is excluded from the 27-point case, as this benchmark is not supported. Reported results are averages over five runs. For our tests, we set the maximum number of iterations to $100$ and the relative residual tolerance to $10^{-16}$. Since our focus is on the cost per iteration rather than on convergence properties, this setup ensures that each implementation performs exactly 100 iterations in all scenarios.
@@
Figure 9 breaks down the dynamic energy consumption per DOF for the CG computation, with GPU and CPU contributions shown as colored segments in each bar (see legend). Under weak scalability, all implementations exhibit stable per-DOF energy consumption as the problem size increases. In other words, the energy required per DOF remains nearly constant when the number of processes grows proportionally to the global problem size, indicating that the additional computational resources are effectively amortized and that none of the implementations introduces significant energy overhead as the scale increases. The energy profile of the un-preconditioned CG solver follows the same relative trends observed for the SpMV operation in 5.1. This is expected, since SpMV is by far the dominant kernel within each iteration, and the improvements achieved by BootCMatchGX over Ginkgo in the standalone SpMV directly carry over to the full solver. In absolute terms, however, the dynamic energy per DOF in CG is about two orders of magnitude higher than in SpMV, reflecting the cumulative cost of performing many iterations rather than a single matrix–vector product. Although additional vector operations are executed in each iteration, their contribution does not alter the overall picture, which remains dictated by the efficiency of the SpMV kernel. This explains why BootCMatchGX (and, in the 7-point stencil case, NVIDIA AmgX) consistently achieve significantly lower dynamic energy consumption per iteration and per DOF compared to Ginkgo, as confirmed by the results in Figures 8 and 9. In summary, as expected, the comparative trends observed in CG largely reflect those already highlighted for SpMV, underlining the central role of the sparse matrix–vector product in defining the solver’s energy efficiency.
@@
5.3 Preconditioned Conjugate Gradient solver
We evaluate the PCG implementations of NVIDIA AmgX and BootCMatchGX under strong and weak scalability conditions. For fairness, AmgX is configured with the matching-based aggregation preconditioner, using aggregates of size 8, as in BootCMatchGX. Both libraries use default settings for the AMG hierarchy (levels and coarsest size) and apply the same smoother (4 $\ell_{1}$-Jacobi iterations) in the V-cycle. For the single-GPU case, the problem size is set to $370^{3}$ DOFs (7-point stencil). All performance measurements reported in the figures represent averages over five independent runs. For these tests, iterations stop at a relative residual tolerance of $10^{-6}$, reflecting realistic simulation settings.
@@ 5.3 Preconditioned Conjugate Gradient solver
Finally, Figure 16 shows the GPU power peaks recorded during the PCG computation. Under weak scalability, BootCMatchGX consistently exhibits lower peaks than NVIDIA AmgX, suggesting that it manages GPU resources more efficiently by maintaining a steady workload and avoiding sudden spikes in power demand. Under strong scalability, power peaks decrease as the number of GPUs increases, reflecting the reduced workload per GPU. However, starting from 8 GPUs, NVIDIA AmgX exhibits lower peaks than BootCMatchGX. This behavior is reflected in the higher per-iteration energy consumption of BootCMatchGX, as shown in Figure 15.

### v2

@@ Abstract
We investigate the energy efficiency of a library designed for parallel computations with sparse matrices. The library leverages high-performance, energy-efficient Graphics Processing Unit (GPU) accelerators to enable large-scale scientific applications. Our primary development objective was to maximize parallel performance and scalability in solving sparse linear systems whose dimensions far exceed the memory capacity of a single node. To this end, we devised methods that expose a high degree of parallelism while optimizing algorithmic implementations for efficient multi-GPU usage. Previous work has already demonstrated the library’s performance efficiency on large-scale systems comprising thousands of NVIDIA GPUs, achieving improvements over state-of-the-art solutions. In this paper, we extend those results by providing energy profiles that address the growing sustainability requirements of modern HPC platforms. We present our methodology and tools for accurate runtime energy measurements of the library’s core components and discuss the findings. Our results confirm that optimizing GPU computations and minimizing data movement across memory and computing nodes reduces both time-to-solution and energy consumption. Moreover, we show that the library delivers substantial advantages over comparable software frameworks on standard benchmarks.
@@ 4.1 CPU and GPU Power
powerMonitor has been adapted from GPowerU to better meet our specific requirements, providing a simple, yet effective, solution for monitoring the power consumption of all GPUs on a compute node. Indeed, it can be used either as an external tool or integrated directly into the application code. In our experiments, we used it as an external tool. In this case, it must be launched before the application starts and stopped once execution is complete. The complete workflow for CPU and GPU power monitoring is shown in Figure 1. powerMonitor creates an output file for each device found on the node, containing a sequence of power samplings along with the corresponding timestamps. These samples can be used to reconstruct the power timeline of each device during the application’s execution. Figure 2 shows an example of a power–time curve. We use a sampling frequency that is higher than the actual NVML reporting resolution on A30 hardware. While this does not increase the intrinsic temporal resolution of the measurement or recover sub-second power variations, it provides a denser set of overlapping estimates of the same averaged signal. This higher sampling density yields a smoother power–time curve and reduces the discretization error when integrating power to estimate energy consumption.
Sampling inaccuracies may affect results reproducibility, particularly for workloads characterized by high-frequency GPU operations that generate short-lived power spikes. In our case study on A30 hardware, power measurements collected via the NVML API are averaged over a 1-second interval. Although this temporal aggregation enhances measurement stability, it may smooth transient fluctuations and partially mask peak power events. In large-scale environments, concurrently polling hundreds of devices can introduce non-negligible software overhead and latency. We highlight that the proposed measurement methodology is independent of the number of cluster nodes used for the test and depends solely on the number of devices per node, as LIKWID and powerMonitor operate locally without requiring internode communication. Nevertheless, in large clusters, the accumulation of small per-device measurement errors may limit the precision of energy-efficiency comparisons across different job executions. In our experiments, power sampling is constrained by NVML’s temporal resolution, and results are averaged over five independent runs to mitigate variability. Since the objective of this study is to provide a comparative assessment of state-of-the-art libraries rather than an exact quantification of absolute energy consumption, the analysis emphasizes relative energy usage in order to identify the most energy-efficient implementation.
@@ 5 Results and Comparisons
In addition to Poisson-related matrices, we consider a set of five symmetric positive definite sparse matrices selected from the SuiteSparse Matrix Collection [35]. The matrices are chosen among SPD problems that do not necessarily arise from the discretization of diffusion-dominated equations on structured grids, thereby providing test cases with diverse sparsity patterns and numerical features that often pose significant challenges for Krylov solvers coupled with AMG preconditioners. Key characteristics of the selected matrices are summarized in Table 1. Since these problems are defined at a fixed size, the analysis is restricted to strong scaling experiments. Although the selected matrices exhibit relatively large dimensions compared to other entries in the collection, their overall size makes it reasonable to limit the evaluation to a small number of accelerators. Accordingly, experiments are performed using up to 4 GPUs.
Table 1: Key characteristics of the selected SPD matrices from the SuiteSparse Matrix Collection.
Matrix Rows Nonzeros Avg. nnz/row
G3_circuit 1 585 478 7 660 826 4.8
af_shell8 504 855 17 579 155 34.8
boneS10 914 898 40 878 708 44.7
ecology2 999 999 4 995 991 5.0
parabolic_fem 525 825 3 674 625 7.0
5.1 Results on Poisson-related benchmarks
5.1.1 SpMV Operation
We begin by analyzing the performance behavior of the SpMV operation under both strong and weak scalability scenarios.For the single-GPU case, the problem size is set to $405^{3}$ DOFs for the 7-point stencil and $260^{3}$ DOFs for the 27-point stencil, corresponding to the largest dimensions that fully saturate the available GPU memory. All performance measurements reported in the figures represent averages over five independent runs, with each run consisting of 100 repetitions of the SpMV computation. All plots include 95% confidence intervals, shown as black vertical lines.
@@ 5.1.1 SpMV Operation
Figure 6 presents the dynamic energy consumption per DOFs. As anticipated from previous results, BootCMatchGX demonstrates significantly higher energy efficiency than Ginkgo—approximately twice as efficient. In both implementations, energy efficiency remains nearly constant with increasing numbers of GPUs under weak scalability, confirming the good scalability of both solutions. Under strong scalability, as expected, efficiency declines as the number of GPUs increases, due to the reduction of the computational workload per GPU.
Finally, Tables 2 and 3 compare static and dynamic energy consumption, expressing the latter as a percentage of the static energy. Although BootCMatchGX shows a higher dynamic energy percentage, this should be interpreted in light of its design strategy. In weak scalability scenarios, the decomposition of static and dynamic energy provides insight into sustainability trade-offs. Since static energy is time-dependent, reductions in communication overhead and idle phases — achieved through improved data locality and communication-aware design — directly limit its contribution. Therefore, the higher dynamic energy share observed for BootCMatchGX likely reflects a larger fraction of energy devoted to useful computation rather than baseline dissipation. When considered together with its shorter runtimes and lower total dynamic energy consumption, this behavior supports the view that BootCMatchGX achieves more effective GPU utilization, with a lower overall energy consumption.
Table 2: Static vs. dynamic energy consumption, for the 7-points stencil matrix with $405^{3}$ DOFs under weak scalability. Columns report GPU, CPU, and total dynamic energy consumption expressed as a percentage of the static energy.
GPU % CPU % total %
#GPUs library
1 BCMGX 47.29 16.12 42.25
Ginkgo 39.77 10.45 36.38
2 BCMGX 63.94 17.61 52.94
Ginkgo 39.52 21.33 37.38
4 BCMGX 61.75 18.83 51.99
Ginkgo 42.03 24.44 39.86
8 BCMGX 61.91 19.00 52.15
Ginkgo 38.54 22.04 36.77
16 BCMGX 51.74 20.61 45.96
Ginkgo 38.99 21.42 37.03
32 BCMGX 60.05 20.41 51.89
Ginkgo 43.06 20.42 40.23
64 BCMGX 51.71 19.82 45.88
Ginkgo 40.78 20.08 38.28
Table 3: Static vs. dynamic energy consumption, for the 27-points stencil matrix with a total of $260^{3}$ DOFs under weak scalability. Columns report GPU, CPU, and total dynamic energy consumption expressed as a percentage of the static energy.
GPU % CPU % total %
#GPUs library
1 BCMGX 36.75 11.81 33.16
Ginkgo 28.95 16.58 28.04
2 BCMGX 45.68 18.81 41.12
Ginkgo 35.58 28.53 34.88
4 BCMGX 46.27 18.08 41.56
Ginkgo 37.23 20.14 35.61
8 BCMGX 47.12 18.68 42.23
Ginkgo 34.73 19.31 33.37
16 BCMGX 40.56 20.28 37.66
Ginkgo 39.44 17.52 37.15
32 BCMGX 41.20 19.04 38.03
Ginkgo 36.77 18.45 35.05
64 BCMGX 42.25 18.56 38.79
Ginkgo 37.08 19.10 35.38
@@
5.1.2 Un-preconditioned Conjugate Gradient solver
@@
For the single-GPU case, the problem size is set to $408^{3}$ DOFs with the 7-point stencil and $265^{3}$ DOFs with the 27-point stencil; AmgX is excluded from the 27-point case, as this benchmark is not supported. Reported results are averages over five runs and include 95% confidence intervals, shown as black vertical lines. For some configurations (1, 32, and 64 GPUs) of the 7-point stencil in the weak scaling settings, we ran also the t-test. The result of the t-test confirms that the null hypotheses (the AmgX, BootCMatchGX, and Ginkgo samples have the same mean) is always rejected. As a consequence, we are confident that the differences are not due to noise in the measures. For our tests, we set the maximum number of iterations to $100$ and the relative residual tolerance to $10^{-16}$. Since our focus is on the cost per iteration rather than on convergence properties, this setup ensures that each implementation performs exactly 100 iterations in all scenarios.
@@
Figure 9 breaks down the dynamic energy consumption per DOF for the CG computation, with GPU and CPU contributions shown as colored segments in each bar (see legend). Under weak scalability, all implementations exhibit stable per-DOF energy consumption as the problem size increases. In other words, the energy required per DOF remains nearly constant when the number of processes grows proportionally to the global problem size, indicating that the additional computational resources are effectively amortized and that none of the implementations introduces significant energy overhead as the scale increases. The energy profile of the un-preconditioned CG solver follows the same relative trends observed for the SpMV operation in 5.1.1. This is expected, since SpMV is by far the dominant kernel within each iteration, and the improvements achieved by BootCMatchGX over Ginkgo in the standalone SpMV directly carry over to the full solver. In absolute terms, however, the dynamic energy per DOF in CG is about two orders of magnitude higher than in SpMV, reflecting the cumulative cost of performing many iterations rather than a single matrix–vector product. Although additional vector operations are executed in each iteration, their contribution does not alter the overall picture, which remains dictated by the efficiency of the SpMV kernel. This explains why BootCMatchGX (and, in the 7-point stencil case, NVIDIA AmgX) consistently achieve significantly lower dynamic energy consumption per iteration and per DOF compared to Ginkgo, as confirmed by the results in Figures 8 and 9. In summary, as expected, the comparative trends observed in CG largely reflect those already highlighted for SpMV, underlining the central role of the sparse matrix–vector product in defining the solver’s energy efficiency.
Tables 4 and 5 compare static and dynamic energy consumption, reporting the latter as a percentage of the static energy. Ginkgo consistently exhibits higher dynamic energy percentages than BootCMatchGX, and its dynamic energy consumption and runtimes are also greater, indicating less efficient GPU utilization and, consequently, higher total energy consumption. AmgX shows higher dynamic energy percentages than both Ginkgo and BootCMatchGX; however, its dynamic energy consumption and runtime are lower than Ginkgo’s, leading to lower overall energy consumption as a result of more effective GPU utilization. Although BootCMatchGX and AmgX achieve comparable runtimes, BootCMatchGX attains lower dynamic energy consumption—and therefore lower total energy consumption—indicating more efficient GPU resource utilization associated with a lower average power draw.
Table 4: Static vs. dynamic energy consumption, for the 7-points stencil matrix with a total of $408^{3}$ DOFs under weak scalability. Columns report GPU, CPU, and total dynamic energy consumption expressed as a percentage of the static energy.
GPU % CPU % total %
#GPUs library
1 AMGX 217.33 12.40 105.53
BCMGX 81.26 12.44 59.17
Ginkgo 111.99 11.07 96.67
2 AMGX 186.69 23.56 105.86
BCMGX 69.32 22.96 55.77
Ginkgo 84.84 23.44 77.59
4 AMGX 172.72 19.62 99.63
BCMGX 71.69 16.71 55.60
Ginkgo 114.80 14.98 99.56
8 AMGX 170.74 19.83 99.01
BCMGX 63.09 16.93 50.67
Ginkgo 94.92 21.31 85.37
16 AMGX 166.05 16.61 95.87
BCMGX 62.91 18.87 51.18
Ginkgo 100.67 17.26 89.14
32 AMGX 169.68 17.20 96.55
BCMGX 71.43 17.84 55.81
Ginkgo 104.61 19.48 92.11
64 AMGX 161.78 18.45 94.16
BCMGX 63.73 18.15 51.49
Ginkgo 102.88 18.91 90.60
Table 5: Static vs. dynamic energy consumption, for the 27-points stencil matrix with a total of $265^{3}$ DOFs under weak scalability. Columns report GPU, CPU, and total dynamic energy consumption expressed as a percentage of the static energy.
GPU % CPU % total %
#GPUs library
1 BCMGX 47.04 9.60 40.27
Ginkgo 77.70 14.69 71.42
2 BCMGX 46.94 17.01 41.38
Ginkgo 84.94 23.38 78.34
4 BCMGX 40.38 19.42 36.82
Ginkgo 82.22 11.40 74.74
8 BCMGX 47.12 19.67 42.05
Ginkgo 89.56 19.81 81.61
16 BCMGX 42.46 19.65 38.53
Ginkgo 83.12 20.80 76.59
32 BCMGX 40.76 19.43 37.15
Ginkgo 79.44 17.69 73.35
64 BCMGX 41.71 17.89 37.77
Ginkgo 78.55 18.49 72.59
@@
5.1.3 Preconditioned Conjugate Gradient solver
We evaluate the PCG implementations of NVIDIA AmgX and BootCMatchGX under strong and weak scalability conditions. For fairness, AmgX is configured with the matching-based aggregation preconditioner, using aggregates of size 8, as in BootCMatchGX. Both libraries use default settings for the AMG hierarchy (levels and coarsest size) and apply the same smoother (4 $\ell_{1}$-Jacobi iterations) in the V-cycle. For the single-GPU case, the problem size is set to $370^{3}$ DOFs (7-point stencil). All performance measurements reported in the figures represent averages over five independent runs and include 95% confidence intervals, shown as black vertical lines. For these tests, iterations stop at a relative residual tolerance of $10^{-6}$, reflecting realistic simulation settings.
@@ 5.1.3 Preconditioned Conjugate Gradient solver
Figure 16 shows the GPU power peaks recorded during the PCG computation. Under weak scalability, BootCMatchGX consistently exhibits lower peaks than NVIDIA AmgX, suggesting that it manages GPU resources more efficiently by maintaining a steady workload and avoiding sudden spikes in power demand. Under strong scalability, power peaks decrease as the number of GPUs increases, reflecting the reduced workload per GPU. However, starting from 8 GPUs, NVIDIA AmgX exhibits lower peaks than BootCMatchGX. This behavior is reflected in the higher per-iteration energy consumption of BootCMatchGX, as shown in Figure 15.
Finally, Table 6 compares static and dynamic energy consumption, expressing the latter as a percentage of the former. AmgX exhibits higher dynamic energy percentages than BootCMatchGX, together with greater dynamic energy consumption and longer runtimes, reflecting less efficient GPU utilization and resulting in higher total energy consumption.
Table 6: Static vs. dynamic energy consumption, for the 7-points stencil matrix with a total of $370^{3}$ DOFs under weak scalability. Columns report GPU, CPU, and total dynamic energy consumption expressed as a percentage of the static energy.
GPU % CPU % total %
#GPUs library
1 AMGX 262.39 18.40 106.57
BootCMatchGX 155.01 10.11 80.18
2 AMGX 294.45 25.88 117.24
BootCMatchGX 172.49 9.58 82.48
4 AMGX 301.32 17.68 112.13
BootCMatchGX 226.17 18.02 97.38
8 AMGX 309.09 14.02 110.24
BootCMatchGX 237.75 18.66 99.85
16 AMGX 329.15 21.06 117.17
BootCMatchGX 218.67 19.35 92.89
32 AMGX 327.52 19.15 115.29
BootCMatchGX 238.84 18.56 96.01
64 AMGX 323.88 18.86 113.92
BootCMatchGX 241.09 19.93 96.36
@@
5.2 Results on Matrices from SuiteSparse
For space reasons, we restrict the discussion to the main results obtained on matrices from the SuiteSparse Collection, summarizing execution times and dynamic energy consumption (GPU, CPU, and total), as well as peak GPU power, for both the SpMV operation and the CG solver. We do not report results obtained using the AMG preconditioners available in the libraries, since AMG methods are not the most appropriate preconditioners for general SPD matrices that do not arise from diffusion-dominated problems, such as those stemming from structural mechanics or optimization applications. Moreover, a detailed analysis of the convergence behavior of AMG preconditioners on general matrices is beyond the scope of this paper.
Table 7: Results with the SpMV Operation on the matrices of Table 1.
time (s) GPU
dynamic energy (J)
CPU
dynamic energy (J)
Total
dynamic energy (J)
GPU power peak (W)
# GPUs matrix library
1 G3_circuit BCMGX 0.0002 0.1637 0.0028 0.1665 82
Ginkgo 0.0004 0.2011 0.0069 0.2080 64
af_shell8 BCMGX 0.0004 0.1907 0.0029 0.1936 100
Ginkgo 0.0006 0.5047 0.0078 0.5125 100
boneS10 BCMGX 0.0012 0.6754 0.0115 0.6869 105
Ginkgo 0.0012 1.1370 0.0170 1.1540 197
ecology2 BCMGX 0.0001 0.0557 0.0023 0.0579 63
Ginkgo 0.0003 0.1363 0.0042 0.1405 64
parabolic_fem BCMGX 0.0001 0.1333 0.0021 0.1354 40
Ginkgo 0.0002 0.0637 0.0024 0.0662 59
2 G3_circuit BCMGX 0.0002 0.2180 0.0096 0.2275 61
Ginkgo 0.0013 0.4741 0.0620 0.5361 46
af_shell8 BCMGX 0.0002 0.2896 0.0070 0.2965 76
Ginkgo 0.0005 0.8035 0.0175 0.8210 82
boneS10 BCMGX 0.0007 0.7326 0.0236 0.7562 102
Ginkgo 0.0009 2.0074 0.0285 2.0359 179
ecology2 BCMGX 0.0001 0.1178 0.0040 0.1218 48
Ginkgo 0.0002 0.2628 0.0104 0.2733 64
parabolic_fem BCMGX 0.0004 0.1641 0.0183 0.1824 42
Ginkgo 0.0018 0.2526 0.1061 0.3587 41
4 G3_circuit BCMGX 0.0004 0.3660 0.0311 0.3971 43
Ginkgo 0.0015 1.1603 0.1143 1.2746 52
af_shell8 BCMGX 0.0001 0.6245 0.0095 0.6340 67
Ginkgo 0.0003 1.6336 0.0218 1.6554 68
boneS10 BCMGX 0.0003 1.1179 0.0265 1.1444 93
Ginkgo 0.0007 3.7145 0.0479 3.7624 146
ecology2 BCMGX 0.0001 0.3321 0.0066 0.3387 45
Ginkgo 0.0002 0.7931 0.0113 0.8044 74
parabolic_fem BCMGX 0.0007 0.4641 0.0695 0.5336 40
Ginkgo 0.0028 0.9583 0.2921 1.2504 58
Table 7 summarizes the SpMV performance results for BootCMatchGX (BCMGX) and Ginkgo. As expected, the scalability trend strongly depends on the sparsity pattern and average number of nonzeros per row. Matrices with a more regular structure and higher average connectivity, such as boneS10 and af_shell8, exhibit better scalability, whereas matrices characterized by irregular sparsity patterns, such as G3_\_circuit, show an increase in execution time for increasing number of GPUs due to increased communication and load imbalance effects. The parabolic_\_fem matrix exhibits a regular sparsity pattern typical of FEM discretizations, however its SpMV performance degrades as the number of GPUs increases, due to the presence of non-zeros far from the diagonal, which leads to increased inter-process communication that negatively impacts scalability. The ecology2 matrix exhibits a moderately irregular sparsity pattern with most non-zeros located close to the diagonal, resulting in limited but non-negligible inter-process communication in multi-GPU SpMV. The energy analysis closely follows the observed runtime behavior. BootCMatchGX exhibits low energy consumption across all configurations except one, the single-GPU parabolic_fem case. Its performance is consistent with shorter runtimes and more effective utilization of GPU resources, as indicated by lower power peaks and average power. The only exception is characterized by higher average power, which leads to greater energy consumption despite a shorter runtime compared to Ginkgo. Regarding GPU power peaks, the results indicate a moderate decrease as the number of GPUs grows, consistently with the reduced workload per device. Across all test cases, BootCMatchGX exhibits lower or comparable power peaks than Ginkgo, suggesting a more stable utilization of GPU resources and a reduced incidence of short-lived power spikes. Overall, the comparison highlights that BootCMatchGX generally outperforms Ginkgo for the SpMV operation also on SuiteSparse matrices, achieving lower execution times, reduced dynamic energy consumption, and more favorable power profiles. These results confirm that the optimizations adopted in BootCMatchGX translate into tangible benefits not only in performance, but also in energy efficiency across a wide range of sparse matrix characteristics.
Table 8: Results with the CG Solver on the matrices of Table 1.
runtime (s) GPU
dynamic energy (J)
CPU
dynamic energy (J)
dynamic energy (J) GPU power peak (W)
# GPUs matrix library
1 G3_circuit AMGX 0.0686 18.7061 1.1757 19.8818 88
BCMGX 0.0588 8.4993 0.8437 9.3430 82
Ginkgo 0.0804 61.5499 0.7746 62.3245 143
af_shell8 AMGX 0.0660 30.6657 1.0121 31.6779 91
BCMGX 0.0643 12.3762 1.0650 13.4412 83
Ginkgo 0.0701 75.3592 1.2008 76.5601 163
boneS10 AMGX 0.1199 77.2129 1.6811 78.8940 100
BCMGX 0.1478 61.1318 2.1697 63.3015 98
Ginkgo 0.1527 166.6234 0.8179 167.4413 175
ecology2 AMGX 0.0503 12.0110 0.5640 12.5750 70
BCMGX 0.0434 2.6365 0.6477 3.2843 64
Ginkgo 0.0612 53.7802 0.6177 54.3979 122
parabolic_fem AMGX 0.0373 13.0647 0.4709 13.5357 63
BCMGX 0.0330 3.6301 0.5339 4.1640 63
Ginkgo 0.0490 24.3215 0.7923 25.1138 85
2 G3_circuit AMGX 0.0674 28.3306 2.6856 31.0162 58
BCMGX 0.0559 12.4181 2.7095 15.1275 63
Ginkgo 0.1671 60.6846 6.1768 66.8614 52
af_shell8 AMGX 0.0501 40.8106 2.6652 43.4758 68
BCMGX 0.0470 17.9323 2.2262 20.1585 76
Ginkgo 0.0663 135.9255 1.8674 137.7929 121
boneS10 AMGX 0.0755 141.1463 3.7273 144.8736 91
BCMGX 0.0943 52.7114 3.6241 56.3355 90
Ginkgo 0.1132 260.4663 4.0540 264.5203 158
ecology2 AMGX 0.0424 11.2354 2.0118 13.2473 56
BCMGX 0.0341 5.4043 1.2162 6.6206 44
Ginkgo 0.0533 57.3245 1.4502 58.7747 74
parabolic_fem AMGX 0.0736 14.4014 3.6070 18.0084 44
BCMGX 0.0670 7.2354 2.7290 9.9644 42
Ginkgo 0.2285 29.8936 11.4415 41.3350 44
4 G3_circuit AMGX 0.0707 53.4715 5.0398 58.5113 45
BCMGX 0.0645 19.6857 4.7947 24.4804 46
Ginkgo 0.2034 83.7765 17.5171 101.2936 59
af_shell8 AMGX 0.0405 92.8779 3.3988 96.2767 57
BCMGX 0.0343 35.9482 2.1871 38.1353 59
Ginkgo 0.0608 164.4908 4.0691 168.5599 83
boneS10 AMGX 0.0543 194.7348 5.4353 200.1701 76
BCMGX 0.0611 75.3093 4.2240 79.5334 78
Ginkgo 0.0832 454.3023 4.9802 459.2825 134
ecology2 AMGX 0.0375 30.6689 2.4062 33.0751 44
BCMGX 0.0290 10.6009 1.6763 12.2772 43
Ginkgo 0.0510 75.3333 2.7597 78.0929 69
parabolic_fem AMGX 0.0946 28.1975 9.7096 37.9071 39
BCMGX 0.0894 19.5766 7.6658 27.2424 40
Ginkgo 0.3167 66.5021 25.2880 91.7900 52
The results reported in Table 8 show trends that are fully consistent with those observed for the SpMV operation in Table 7. This behavior is expected, since SpMV represents the dominant computational kernel within each CG iteration, accounting for the largest fraction of both execution time and dynamic energy consumption. As the number of GPUs increases, the CG execution time follows the same scaling patterns observed for SpMV across the different matrices. In particular, matrices exhibiting favorable SpMV scalability, such as boneS10 and af_\_shell8, also show improved CG performance, whereas matrices characterized by irregular sparsity patterns or large off-diagonal contributions, such as G3_\_circuit and parabolic_\_fem, exhibit a degradation in CG scalability due to increased communication overheads. The ecology2 matrix again represents an intermediate case, with moderate scalability limitations reflecting its sparsity structure. The energy consumption trends closely mirror the runtime behavior. Since the energy-to-solution of CG is largely dictated by the cumulative cost of repeated SpMV operations, reductions (or increases) in execution time directly translate into corresponding changes in dynamic energy consumption. Similarly, GPU power peaks show the same qualitative behavior observed for SpMV, decreasing with the number of GPUs as the per-device workload is reduced. Overall, the CG results confirm that improvements achieved at the SpMV level carry over directly to the full solver. In this respect, BootCMatchGX maintains the advantages observed in SpMV also for CG, achieving lower execution times, reduced dynamic energy consumption, and more stable power profiles compared to Ginkgo and AmgX across all tested matrices.